# LayerCollapse: Adaptive compression of neural networks

### v1

@@ 3 Methodology
where, $n_{in}$, $n_{hidden}$, and $n_{out}$ denote the input size, hidden size, and output size of a 2-layer MLP, respectively.
@@
An interesting aspect of this formula is that the gain can sometimes be negative. Positive gain happens when $n_{hidden}$ is larger than the product of $n_{in}$ and $n_{out}$ divided by their sum. This means that MLPs designed with a “bottleneck” structure (where $n_{hidden}$ is relatively small) may not see much benefit from our layer collapsing method. On the other hand, most modern MLP designs, which tend to expand the layer size (“widening” structure), are likely to experience significant positive gains from our method.
3.1 Adaptation for Convolutional Layers
@@
Convolutional layers play a pivotal role in numerous neural network architectures. Essentially convolution can be formulated as sparse fully connected layers with some redundancies. Their nature as linear transforms permits representation via circulant matrices. A circulant matrix is a matrix where the row vectors have the same elements and each row vector is rotated one element to the right relative to the preceding row vector. An interesting property of these matrices is that their product also forms a circulant matrix, resulting in another convolution layer. The kernel size of the resulting convolution layer, denoted as $k$, is calculated by $k=k_{1}+k_{2}-1$ where, $k_{1}$ and $k_{2}$ are the kernel sizes of the original convolution layers. The compression gain can be calculated by:
$Gain_{conv}=1-\frac{k_{1}^{2}\times c_{in}\times c_{hidden}+k_{2}^{2}\times c_{hidden}\times c_{out}}{(k_{1}+k_{2}-1)^{2}\times c_{in}\times c_{out}}$ (8)
@@
where, $c_{in}$, $c_{hidden}$, and $k_{1}$ correspond to the input channels, output channels, and the kernel size of the initial convolution layer, respectively. Meanwhile, $c_{hidden}$, $c_{out}$ and $k_{2}$ are the parameters of the subsequent convolution layer. However, an essential note when applying LayerCollapse to common CNN architectures, such as ResNet, is that the gain often turns out to be negative. To achieve a positive gain, it is imperative for $c_{hidden}$ to be significantly larger than both $c_{in}$ and $c_{out}$. This observation highlights the need for careful configuration when attempting to collapse layers in such architectures. Overall, this framework offers a principled approach to explore the potential benefits of LayerCollapse on convolution layers, illuminating the optimization potential for convolutional neural networks.
3.2 Prior Distribution Regularization
@@
We extend the use of the parametric ReLU (PReLU) function to develop a regularization method based on the maximum a posteriori (MAP) framework. Aligning with Occam’s razor, which favors simpler explanations, we propose to adopt a normal distribution with a mean of 1 as the prior distribution for $\alpha$. This setting effectively limits the two layers to a linear space, reducing their functional complexity, as illustrated in Figure 1. This choice is backed by empirical results given in section 4. MAP framework suggests adding negative log likelihood of the prior distribution of parameters to the loss function. This is referred to as the prior loss or regularization loss. Following this principal the added regularization term is:
@@
where $lc$ is the regularization constant.
@@
1:function Collapse($W_{1},b_{1},W_{2},b_{2},\gamma,\beta,\bar{\mu},\bar{\sigma},\alpha,\tau$)
2:     if $\alpha\leq\tau$ then
3:         $s_{BN}\leftarrow\frac{\gamma}{\sqrt{\bar{\sigma}}}$
4:         $W_{\text{new}}\leftarrow s_{BN}\times W_{2}\cdot W_{1}$
5:         $b_{\text{new}}\leftarrow W_{2}\cdot(s_{BN}\times(b_{1}-\bar{\mu})+\beta)+b_{2}$
6:         return $W_{\text{new}},b_{\text{new}}$
7:     else
8:         return un-collapsible
9:     end if
10:end function
@@ 4 Experiments
In this section we evaluate our method on both post training compression, and compression aware training with our regularization. Initially, we explore the standalone capability of LayerCollapse in post training compression of streamlined vision models including vision transformer, mlp-mixer and VGG. We then compare LayerCollapse with Knowledge Distillation on the cost of training and achievable results on a target network. Subsequently, we demonstrate the effectiveness of compression aware training in an ablation study on mid-sized datasets: TinyImageNet [22], and CIFAR100 [21]. Further extending the scope of our investigation, we exhibit preliminary results on the BERT [2] language model to provide a proof-of-concept for the generalizability of LayerCollapse in domains beyond computer vision, such as natural language processing (NLP). These experiments aim to demonstrate LayerCollapse’s versatility and its potential to enhance efficiency and performance in various machine learning contexts.
Model Collapsed Layers No. of Params (M) Top1 Acc. (%) Top5 Acc. (%) MACs (G)
ViT T/16 0 5.72 75.35 93 1.08
1 5.46 75.79 93 1.02
2 5.20 75.35 92 0.97
3 4.94 74.42 92 0.91
ViT S/16 0 22.05 81.38 96 4.24
1 21.02 80.96 96 4.04
2 19.98 79.74 95 3.84
ViT B/16 0 86.57 81.43 96 16.85
1 82.44 82.17 96 16.04
2 78.3 80.31 95 15.23
ViT L/16 0 304.33 84.80 98 59.66
2 289.64 84.22 97 56.77
Mixer B/16 0 59.88 76.47 92 12.61
2 55.64 76.11 92 11.71
4 51.39 75.67 91 10.81
VGG19 0 143.67 74.21 92 19.65
2 45.11 71.22 88 19.55
VGG16 0 138.36 73.36 92 15.48
2 39.80 67.16 88 15.39
VGG13 0 133.05 71.58 90 11.32
2 34.49 67.21 88 11.22
VGG11 0 132.86 70.37 90 7.62
2 34.31 66.36 88 7.52
Table 2: Post training compression results, showcasing accuracy and compression attained via LayerCollapse across different layers in relevant architectures.
4.1 Compression
Our compression experiments are performed on ImageNet-1k classification task [30]. Unlike pruning [24] and quantization [15] compression methods which require finetuning on the compressed method to recover accuracy, LayerCollapse utilizes finetuning with the regularizer to ensure accuracy. Afterwards the compression phase will endure no accuracy loss. We first assess the stand-alone compression efficacy of LayerCollapse. This is followed by a comparative analysis of LayerCollapse and Knowledge Distillation, focusing on training efficiency and resource utilization.
4.1.1 Stand-Alone Compression
Dataset. We employ the ImageNet-1k [30] dataset for both fine-tuning and validation processes. It comprises over 1.2 million images for training and 50,000 for validation.
Implementation. We adapt pre-trained models from the TIMM library [39] for MLP-Mixer [35] and ViT [7], alongside VGG variants from torchvision [28]. ViT and Mixer architectures use GeLU activations in their original implementation. In target layers, these are converted to ReLU to align with the requirements of LayerCollapse. The fine-tuning follows a consistent procedure across models, using Stochastic Gradient Descent (SGD) with momentum and a learning rate of $5e^{-4}$ as the optimizer. The loss function is comprised of three parts, cross entropy between labels and result, KL divergence between the output and uncollapsed models output and the LayerCollapse regularization term (9) with $lc=0.2$. This process is capped at 20 epochs. Each layer for all ViT experiments use 5 epochs. Number of epochs for VGG architectures are given in Table 3. The MAC operations are all calculated by propagating a random $224\times 224$ image through the model. Further details are enumerated in Appendix B.
Sensitivity Analysis. To quantify the robustness of LayerCollapse, we conduct a layer-by-layer sensitivity analysis on the Vision Transformer T/16 model, as shown in Figure 3. This removal of layers provides insight into the impact of incremental compression on the model’s accuracy. Interestingly, we observe an initial increase in accuracy upon the collapse of the first layer, indicating an improvement in generalization compared to the original model.
Model Method Top1 Acc(%) Top5 Acc(%) No. of Parameters (M) Cost Epochs
Initial Final
vgg19 LC 71.22 88 143.67 45.11 11h 5
KD 63.42 85 51h 23
vgg16 LC 70.01 89 138.36 39.80 14h 7
KD 59.63 82 31h 16
vgg13 LC 67.21 88 133.05 34.49 16h 10
KD 58.26 81 45h 28
vgg11 LC 66.36 88 132.86 34.31 20h 20
KD 58.64 81 42h 42
Table 3: A comparative analysis of Layer Collapse (LC) and Knowledge Distillation (KD) methods applied to VGG models.
Discussion. Table 2 illustrates that LayerCollapse achieves a substantial reduction in model size—up to 74% in VGG11—and computational demands with minimal accuracy degradation. Notably, models with heavier classification heads, such as VGG, can benefit significantly from our method. ViT models, the T/16 and B/16 variants, get an increase in accuracy with the first layer collapse. This can be attributed to an increase of generalization in simpler models. In ViT T/16, S/16, and B/16 we get roughly 5 percent reduction in model size per LayerCollapse. In all ViT variants we get a consistent reduction of over 10% in size and complexity with less than a 1% accuracy loss. In the Mixer architecture, 2 types of MLP layers known as MLP-token and MLP Channels exist. MLP-Channel has a significantly bigger size. We found that MLP-tokens can be collapsed in about an epoch of computation whilst MLP-channels require more effort. In the Mixer experiment, we collapsed the layers in pairs to ensure MLP-Channels are affected. In the Mixer architecture we get 14% reduction in size with less than a 1% accuracy degradation. Notice that the Mixer architecture is the only one that gets a linear reduction in MAC operations with each pair collapsed. We used the idea first introduced in Knowledge Distillation to enhance the performance of our finetuning procedure. Using the KL-Divergence loss, we speedup the gradient flow and increase the convergence speed.
4.2 LayerCollapse vs Knowledge Distillation
Setup. Our evaluation framework utilizes batch-normalized VGG variants from the torchvision repository [28], with ImageNet-1K [30] providing a substantial dataset for training and validation. These variants are chosen for their prevalent use in compression studies and their capability to learn without prior training. To facilitate a fair comparison between LayerCollapse and Knowledge Distillation, we adopt the same target architecture for both methods. The rest of the training is set up identical to section 4.1.1 for both LayerCollapse and Knowledge distillation.
Comparison. The models are rigorously compared across several dimensions: training cost, the number of training epochs required, and the accuracy metrics—top-1 and top-5—on the validation set. The computational costs are quantified on an Nvidia RTX A6000 GPU.
Discussion. Referencing Table 3, LayerCollapse consistently achieves over 10% higher accuracy while incurring only 30% of the computational cost compared to Knowledge Distillation. This trend is more pronounced in larger models, which exhibit greater compression potential, and result in up to 15% accuracy gain with 50% of tuning cost. It is noteworthy that for models necessitating pre-training, such as MLP-Mixer and ViT, the absence of pre-training data and the high computational demands often make Knowledge Distillation impractical. LayerCollapse emerges as a unique method for model compression through architectural modification, offering significant cost advantages as demonstrated in Table 3. However, it is important to recognize the inherent limitations of LayerCollapse in terms of the diversity of achievable target model architectures.
4.3 Ablation Study of Regularization
@@
Dataset. Our study utilizes the CIFAR100 [10] and TinyImageNet [22] datasets, which are mid-sized and rich in diversity. The CIFAR100 dataset includes 60,000 images across 100 classes, and TinyImageNet, a subset of the larger ImageNet, comprises 100,000 images across 200 classes.
Implementation. The experiments are conducted using the VGG architecture because of its effective learning capabilities without pre-training and competitive performance in benchmarks. For consistency, all experiments utilize the same setup and hyper parameters. Note that the goal of this study is not to show state of the art performance using our proposed regularization, but to show its effectiveness in reducing over-fitting and final model size. In all experiments, traditional regularizers, such as data augmentation (rotation, random crop, and flipping), dropout[5], and batch normalization[16], have been used. This is to show the added contribution of 9. Training is done using Stochastic Gradient Descent (SGD) with momentum a learning rate scheduler, trained for a total of 400 epochs per experiment. The regularization is applied to 40% of the layers, starting with the latest layers, with the strength of 0.05 across all experiments. extensive experiment details can be found in Appendix B.
@@
4.4 LayerCollapse in NLP
We extended LayerCollapse to the NLP domain. This section presents our initial findings using the BERT-Base model [2], a well-established pre-trained NLP model.
Setup In our evaluation we use the General Language Understanding Evaluation on the Microsoft Research Paraphrase Corpus (GLUE-MRPC)[37][6]. This benchmark assesses a model’s capacity to discern paraphrased sentences, determining if they are equivalent. Further details are found in Appendix B
Discussion We report our results in Table 5. These include accuracy metrics and the number of parameters for each layer of BERT-Base modified by LayerCollapse. The experiment shows a promising 19% reduction in parameters with a minimal accuracy loss of less than 3%. Notably, we exclusively use the data in the GLUE-MRPC task in the LayerCollapse procedure. These preliminary results in the NLP sector suggest the versatility and potential of LayerCollapse across different domains, not limited to computer vision. They point to the method’s adaptability and effectiveness in optimizing model parameters while preserving accuracy without the need for pre-train dataset.
@@ 4.4 LayerCollapse in NLP
Table 5: Layer-wise analysis of accuracy and parameters for Bert-Base via LC
Layers Accuracy (%) Params (M)
0 $85.0$ $109.40$
1 $85.8$ $105.35$
2 $84.1$ $101.20$
3 $84.3$ $97.10$
4 $82.1$ $92.95$
5 $82.3$ $88.80$
6 $82.3$ $84.70$
7 $80.4$ $80.56$
8 $79.4$ $76.40$
5 Conclusion
This paper introduces LayerCollapse, a novel method for compressing neural networks through regularization. LayerCollapse simplifies a network by merging pairs of linear layers, reducing the network’s depth, the number of parameters, thereby, reducing operational requirements. We introduce a compression aware regularizer that encourages the development of shallower network architectures.
Under rigorous evaluation, LayerCollapse demonstrates state-of-the-art performance in compressing Vision Transformers, improving even improving accuracy in some cases. In comparison to Knowledge Distillation, LayerCollapse consistently demonstrated superior efficiency in both computation and final results, achieving over 10% better performance while utilizing 80% fewer resources. LayerCollapse marks a significant paradigm shift in efficient neural network design and bridges the gap between performance and model size.
Supplementary Material

### v3

@@ 3.2 Compression Ratio
where $CR$ is the compression ratio, $n_{in}$, $n_{hidden}$, and $n_{out}$ denote the input size, hidden size, and output size of the 2-layer MLP, respectively.
@@
Noteworthy is that the compression ratio can be negative. Positive compression ratio only happens when $n_{hidden}$ is larger than the product of $n_{in}$ and $n_{out}$ divided by their sum. This means that MLPs designed with a “bottleneck” structure (where $n_{hidden}$ is relatively small) may not see a benefit from LayerCollapse . However, most modern MLP designs tend to expand the layer size (“widening” structure), and are likely to have significant compression ratios.
@@
Convolutional layers play a pivotal role in numerous neural network architectures. Essentially, convolution can be formulated as sparse fully connected layers with some redundancies. Their nature as linear transforms permits representation via circulant matrices. A circulant matrix is a matrix where the row vectors have the same elements and each row vector is rotated one element to the right relative to the preceding row vector. The product of two circulant matrices is also a circulant matrix. So collapsing two convolution layers will result in another convolution layer. The kernel size of the resulting convolution layer, denoted as $k$, is calculated as $k=k_{1}+k_{2}-1$ where, $k_{1}$ and $k_{2}$ are the kernel sizes of the original convolution layers. Compression gain for convolution layers can be calculated as:
$CR_{conv}=1-\frac{k_{1}^{2}\times c_{in}\times c_{hidden}+k_{2}^{2}\times c_{hidden}\times c_{out}}{(k_{1}+k_{2}-1)^{2}\times c_{in}\times c_{out}}$ (8)
@@
where, $c_{in}$, $c_{hidden}$, and $k_{1}$ correspond to the input channels, output channels, and the kernel size of the initial convolution layer, respectively. $c_{hidden}$, $c_{out}$ and $k_{2}$ are the parameters of the subsequent convolution layer. However, a crucial point to consider when applying LayerCollapse to common CNN architectures, such as ResNet, is that the compression ratio is negative. This is due to the fact that a positive compression ratio requires for $c_{hidden}$ to be significantly larger than both $c_{in}$ and $c_{out}$. This observation highlights the need for careful consideration when attempting to collapse layers in such architectures.
3.3 Prior Distribution Analysis of Regularization
@@
We extend the use of the parametric ReLU (PReLU) function to develop a regularization method based on the maximum a posteriori (MAP) framework DeGroot (2005). Aligning with Occam’s razor Schaffer (2015), which favors simpler explanations, we propose to adopt a normal distribution with a mean of 1 as the prior distribution for $\alpha$. This setting effectively limits the two layers to a linear space, reducing their functional complexity. This choice is backed by empirical results given in section 4. An intuition for this choice is presented in figure 1 MAP framework suggests adding negative log likelihood of the prior distribution of parameters to the loss function. This is referred to as the prior loss or regularization loss. Following this principal the added regularization term is:
@@
where $lc$ is the regularization constant. Notice this formulation aligns perfectly to our intuition from equation (4).
@@
  function Collapse($W_{1},b_{1},W_{2},b_{2},\gamma,\beta,\bar{\mu},\bar{\sigma},\alpha,\tau$
     if $\alpha\leq\tau$ then
        $s_{BN}\leftarrow\frac{\gamma}{\sqrt{\bar{\sigma}}}$
        $W_{\text{new}}\leftarrow s_{BN}\times W_{2}\cdot W_{1}$
        $b_{\text{new}}\leftarrow W_{2}\cdot(s_{BN}\times(b_{1}-\bar{\mu})+\beta)+b_{2}$
        return $W_{\text{new}},b_{\text{new}}$
     else
        return un-collapsible
     end if
  end function
Table 2: Training performance on Wikitext-103 dataset. For base variants the modified layer is the last layer i.e. layer 12. In the large variant layers 11, 16, and 32 are selected for compression.
Model Layers Params (M) PPL
GPT-2 Radford et al. (2019) 0 124.4 18.87
GPT-2 + LayerDrop Fan et al. (2019) 1 119.9(-4%) 19.29
GPT-2 + LayerCollapse (ours) 1 120.3(-3%) 19.09
GPT-2 Large Radford et al. (2019) 0 774.0 17.65
GPT-2 Large + LayerDrop Fan et al. (2019) 3 734.7(-5%) 15.61
GPT-2 Large + LayerCollapse (ours) 3 739.6(-4%) 15.17
3.4 Alternative Activation Functions
Thus far, we have focused on the ReLU function as a primary non-linear activation function. ReLU’s simplicity and efficiency have had a significant impact on deep learning since its introduction. Recently, several ReLU variants, including ELU Clevert (2015), GeLU Hendrycks & Gimpel (2016), and SiLU Elfwing et al. (2018), have gained popularity due to their enhanced flexibility.
We propose a general approach to smoothly linearize ReLU-family activations. For functions like ELU, which are linear in the positive region and non-linear in the negative, we incorporate a linear term, $y=x$, into the negative region. This term is blended with the original function using a weighted sum, where the weight parameter is denoted by $\alpha$ as defined in section 3.1. For example, the modified ELU function can be written as:
$\text{ELU}^{\beta}(x)=\begin{cases}x&\text{if }x>0\\ \alpha x+(1-\alpha)\beta(\exp(x)-1)&\text{if }x\leq 0\end{cases}$
where $\beta$ is the ELU parameter. When $\alpha=1$, ELU transitions to a linear function. To ensure $\alpha$ remains within $[0,1]$ during training, we apply a sigmoid to it.
For one-region activations like GeLU and SiLU, which follow the form $f(x)=x\times h(x)$ with $h(x)\rightarrow 1$ as $x\rightarrow\infty$ and $h(x)\rightarrow 0$ as $x\rightarrow-\infty$, we adjust using $\alpha$ in the form $f(x)=x\times(h(x)+\alpha(1-h(x)))$, where $\alpha\in[0,1]$. This leads to parametric versions:
$\text{GeLU}^{\alpha}(x)=x(\Phi(x)+\alpha(1-\Phi(x)))$
where $\Phi(x)$ is the Gaussian cumulative distribution function, and for SiLU:
$\text{SiLU}^{\alpha}(x)=x(\sigma(x)+\alpha(1-\sigma(x)))$
where $\sigma(x)=\frac{e^{x}}{1+e^{x}}$ is the sigmoid function.
@@ 4 Experiments
We conduct a comprehensive evaluation of LayerCollapse ’s accuracy and efficiency. For accuracy, we demonstrate LayerCollapse ’s capability on NLP and vision tasks, showcasing its effectiveness as both a regularizer and a compression method.
Experiments presented in Section 4.1 illustrate LayerCollapse ’s dual role not only as a compression tool for pre-trained networks but also in enhancing model generalization. We compare LayerCollapse against a modified version of LayerDrop, a structured pruning technique, highlighting the distinct advantages of our approach.
Detailed in Section 4.2, we finetune models with LayerCollapse for regularization followed by compression. Additionally, we compare LayerCollapse against Knowledge Distillation, as discussed in Hinton et al. (2015), focusing on the computational resources required for each method, despite targeting the same network architectures.
Finally, we test LayerCollapse as a regularizer by training a VGG network from scratch, presenting two scenarios: using LayerCollapse and not using it, to evaluate its impact on model performance and efficiency.
Table 3: Post-training compression GLUE test results illustrating accuracy and achieved compression. For QQP, MRPC average of F1 and accuracy are reported, for STS-B average of Spearman and Pearson is reported.
Model Layers Params (M) MNLI(m/mm) SST-2 STS-B RTE QNLI QQP MRPC CoLA Average
Bert Base 0 109.1 83.5/83.7 92.7 86.4 65.7 91.0 89.4 85.8 61.0 81.9
4 92.5(-15%) 82.6/82.8 91.2 87.3 61.4 89.8 86.3 88.4 59.6 80.8
Bert Large 0 334.9 86.3/86.1 93.5 87.7 66.4 89.3 89.9 85.8 62.6 82.7
4 309.7(-8%) 86.6/85.5 93.0 89.5 71.5 91.4 89.4 87.5 62.6 83.8
4.1 Regularization Performance of LayerCollapse
4.1.1 Pre-Trained Models
Setup. In our experiments, we utilize the base and large versions of the GPT-2 causal language model Radford et al. (2019), with a fixed context size of 128 tokens. Our layer selection methodology, inspired by Fan et al. (2019), involves initially collapsing each layer, using the algorithm presented in 1, in a one-shot manner, followed by an evaluation to identify layers most tolerant to compression. We evaluate LayerCollapse on the widely used perplexity metric on the test set of the Wikitext-103 dataset Merity et al. (2016).
We conduct training for $3k$ steps, maintaining a batch size of 128 across all experiments. The learning rate scheduling is managed by LiveTune Shabgahi et al. (2023), starting from an initial rate of $0.001$. Optimization is performed using a modified version of the Adam optimizer Kingma & Ba (2014), with comprehensive details about this optimizer available in Appendix B. An exhaustive list of experiment details can be found in Appendix C. The comparison was done with a modified version of LayerDrop Fan et al. (2019). This modification, drawing inspiration from Savarese & Figueiredo (2017), employs gates for selected layers. These gates use a sigmoid function to dynamically alternate between activation and deactivation, enhancing training stability.
Results. As shown in Table 2, LayerCollapse surpasses LayerDrop in terms of final perplexity score. Notably, the large variant of GPT-2 tends to overfit, resulting in higher Perplexity compared to both regularized approaches. However, LayerDrop achieves greater compression gain.
4.1.2 Trained from Scratch.
@@ 4.1.1 Pre-Trained Models
Dataset. Our study utilizes the CIFAR100 Han et al. (2015) and TinyImageNet Le & Yang (2015) datasets, which are mid-sized and rich in diversity. The CIFAR100 dataset includes 60,000 images across 100 classes, and TinyImageNet, a subset of the larger ImageNet, comprises 100,000 images across 200 classes.
Implementation. The experiments are conducted using the VGG architecture because of its effective learning capabilities without pre-training and competitive performance in benchmarks. For consistency, all experiments utilize the same setup and hyper parameters. Note that the goal of this study is not to show state of the art performance using our proposed regularization, but to show its effectiveness in reducing over-fitting and final model size. In all experiments, traditional regularizers, such as data augmentation (rotation, random crop, and flipping), dropoutdo Santos et al. (2021), and batch normalizationIoffe & Szegedy (2015), have been used. This is to show the added contribution of 9. Training is done using Stochastic Gradient Descent (SGD) with momentum a learning rate scheduler, trained for a total of 400 epochs per experiment. The regularization is applied to 40% of the layers, starting with the latest layers, with the strength of 0.05 across all experiments. extensive experiment details can be found in Appendix C.4.
@@
4.2 LayerCollapse for Compression
We apply LayerCollapse to compress targeted layers in pretrained models in both computer vision and NLP tasks. To utilize LayerCollapse as compression for pre-trained models, a few rounds of fine-tuning with the LayerCollapse regularizer are required to ensure the linearization does not significantly impact accuracy. The process involves a few training steps with the regularization loss defined in Equation 9. In all cases, we set the slope threshold for layer collapse at $0.0001$. It’s important to note that we report results directly after layer collapse, without the need for additional fine-tuning.
For architectures such as Bert, ViT, and MLP-Mixer, we substitute GeLU activations in the compressed layers with ReLU, as GeLU is a smoothed ReLU and this change has been found to be insignificant in our tests. Our compression approach focuses on the model’s final layers, driven by the insight that these layers handle more abstract data representations, rendering them more suitable for compression. The compression is carried out one layer at a time. For detailed implementations notes, please refer to Appendix C.
4.2.1 Performance on GLUE
Setup. We assess LayerCollapse ’s compression effectiveness on pre-trained Bert-Large and Bert-Base Devlin et al. (2018) using the GLUE benchmark Wang et al. (2018).The Average column is slightly different than GLUE score since we exclude WNLI as discussed in Devlin et al. (2018).
Results. As illustrated in table 3 LayerCollapse yields $15\%$ compression gain for Bert base with about $1\%$ score reduction on the GLUE benchmark. As evident the larger variant of Bert is prone to overfit thus LayerCollapse regularization improves the score in many tasks. As illustrated LayerCollapse provided better scores than the baseline in Devlin et al. (2018). This is while the resulting model is $8\%$ smaller in terms of parameter count.
Table 5: Post-training compression results, showcasing accuracy and compression attained via LayerCollapse across different layers in relevant architectures.
Model Layers Params (M) MACs (G) Top1 Acc. (%)
ViT T/16 0 5.72 1.08 75.35
3 4.94(-14%) 0.91(-16%) 74.42
ViT S/16 0 22.05 4.24 81.38
2 19.98(-9%) 3.84(-9%) 79.74
ViT B/16 0 86.57 16.85 81.43
2 78.30(-10%) 15.23(-10%) 80.31
ViT L/16 0 304.33 59.66 84.80
2 289.64(-5%) 56.77(-5%) 84.22
Mixer B/16 0 59.88 12.61 76.47
4 51.39(-14%) 10.81(-14%) 75.67
VGG19 0 143.67 19.65 74.21
2 45.11(-67%) 19.55(-1%) 71.22
VGG11 0 132.86 7.62 70.37
2 34.31(-74%) 7.52(-1%) 66.36
4.2.2 Performance on ImageNet-1K
Setup. We apply LayerCollapse across a diverse set of vision models, including CNNs, MLP-Mixers Tolstikhin et al. (2021), and Vision Transformers Dosovitskiy et al. (2020), utilizing the ImageNet-1K benchmark Russakovsky et al. (2015) for evaluation. Pretrained models are sourced from the timm library Wightman (2019). Optimization is performed using SGD with momentum, with a learning rate set to $5e^{-4}$. We employ the Self Distillation method Fan et al. (2019). The loss function is comprised of three parts, cross entropy between labels and result, KL divergence between the output and original model output as the teacher and the LayerCollapse regularization term given in equation (9) with $lc=0.2$. This process is capped at 20 epochs. Details are enumerated in Appendix C.
Sensitivity Analysis. To quantify the robustness of LayerCollapse, we conduct a layer-by-layer sensitivity analysis on the Vision Transformer T/16 model, as shown in Figure 4. This removal of layers provides insight into the impact of incremental compression on the model’s accuracy. We observe an initial increase in accuracy upon the collapse of the first layer, attributed to an improvement in generalization compared to the original model.
Results. Table 5 illustrates that LayerCollapse achieves a substantial reduction in model size-up to $74\%$ in VGG11—with minimal accuracy degradation. Notably, models with heavier classification heads, such as VGG, can benefit significantly from our method. In all ViT variants we get a consistent reduction of over $10\%$ in size and complexity with less than a 1% accuracy loss. In the Mixer architecture, 2 types of MLP layers known as MLP-token and MLP Channels exist. MLP-Channel has a significantly bigger size. We found that MLP-tokens can be collapsed in about an epoch of computation whilst MLP-channels require more effort. In the Mixer experiment, we collapsed the layers in pairs to ensure MLP-Channels are affected. In the Mixer architecture we get $14\%$ reduction in size with less than a $1\%$ accuracy degradation.
4.3 LayerCollapse vs Knowledge Distillation
Setup. Our evaluation framework utilizes batch-normalized VGG variants from the torchvision repository maintainers & contributors (2016). We use the ImageNet-1K Russakovsky et al. (2015) dataset, providing a substantial dataset for training and validation. These variants are chosen for their prevalent use in compression studies. To facilitate a fair comparison between LayerCollapse and Knowledge Distillation, we adopt the same target architecture for both methods. The rest of the training is set up identical to section 4.2 for both LayerCollapse and Knowledge distillation.
Comparison. The models are rigorously compared across several dimensions: training cost, the number of training epochs required, and the accuracy metrics—top-1 and top-5—on the validation set. The computational costs are quantified on one Nvidia RTX A6000 GPU.
Results. Referencing table 6, LayerCollapse consistently achieves over 10% higher accuracy while incurring only 30% of the computational cost compared to Knowledge Distillation. This trend is more pronounced in larger variants, which exhibit greater compression potential. Showing up to 15% accuracy gain with 50% of tuning cost. It is noteworthy that for models necessitating pre-training, such as MLP-Mixer and ViT, the absence of pre-training data and the high computational demands often make Knowledge Distillation impractical. However, we recognize the inherent limitations of LayerCollapse in terms of the diversity of achievable target model architectures. LayerCollapse can only remove fully connected layers while model distillation can change the architecture completely.
5 Discussion
In this study, we presented LayerCollapse , a novel approach for compressing neural networks through regularization techniques that strategically merge linear layer pairs, effectively reducing both network depth and parameter count. This approach specifically encourages the formation of shallower network structures, contributing to computational efficiency. Evaluations show that LayerCollapse achieves better performance in regularization and compression than other structural modifications. Compared to Knowledge Distillation, LayerCollapse delivers over 10% higher performance while using 80% less computational resources. However, the method’s strengths primarily benefit architectures featuring MLPs with widening designs, and it does not substantially improve compression for current CNN models.
Table 6: An analysis of LayerCollapse (LC) and Knowledge Distillation (KD) methods applied to VGG variants.
Model Method Top1 Acc(%) No. of Parameters (M) Cost Epochs
vgg19 LC 71.22 45.11 11h 5
KD 63.42 45.11 51h 23
vgg16 LC 70.01 39.80 14h 7
KD 59.63 39.80 31h 16
vgg13 LC 67.21 34.49 16h 10
KD 58.26 34.49 45h 28
vgg11 LC 66.36 34.31 20h 20
KD 58.64 34.31 42h 42
6 Related Work
In overparameterized neural networks, two major concerns are generalizability and computational requirements. Compression techniques aim to address the second issue by reducing computational demands while maintaining model precision. The predominant methods for model compression include pruning Han et al. (2015), quantization Hubara et al. (2017), and knowledge distillation Hinton et al. (2015), each contributing uniquely to model optimization Gupta & Agrawal (2022). Quantization reduces the number of bits needed to store weights, though it requires hardware support to be effective. Pruning strategies fall into two categories, structured and unstructured. Unstructured pruning reduces storage needs by utilizing sparse representations, but high prune rates and hardware support are necessary. Structured pruning Fan et al. (2019) Lemaire et al. (2019) Fu et al. (2022) changes the model architecture, obviating the need for additional hardware or software support. Knowledge distillation changes the model’s architecture by transferring knowledge from a large model to a smaller one. The problem with knowledge distillation is the immense computational demands of training the target architecture. LayerCollapse , similar to structured pruning and knowledge distillation, modifies the neural network architecture to achieve compression without requiring further software or hardware optimization. Our method compares favorably to LayerDrop Fan et al. (2019) and knowledge distillation Hinton et al. (2015), as shown in Section 4. Pruning and quantization can be iterative or one-shot. Iterative methods compress the model in steps with finetuning, while one-shot methods incorporate regularization during training to enable compression without further finetuning. Quantization-aware training (QAT) Nguyen et al. (2020) and pruning-aware training (PAT) Lemaire et al. (2019)Jiang et al. (2023)Wen et al. (2016) Alvarez & Salzmann (2016) are common practices to mitigate accuracy loss. LayerDrop is a notable PAT that improves training results as a regularizer Fan et al. (2019). In Section 4.1, we evaluate the regularization effect of LayerCollapse against LayerDrop.
Explicit regularization methods, such as Dropout do Santos et al. (2021), Ghiasi et al. (2018), Srivastava et al. (2014), and weight decay, follow the principle that simpler models are more robust Schaffer (2015) and less prone to overfitting. Following the same principle, we adjust the model’s structure with LayerCollapse to prevent overfitting. However, this principle has been shown to be inaccurate in recent over-parameterized models when employed with immense amounts of data.
Previous work on private inference has explored eliminating non-linearities in neural networks for computational efficiency of private inference. DeepReDuce Jha et al. (2021) and Selective Network Linearization (SNL) Cho et al. (2022) adopt regularization strategies for linearization. Unlike SNL, LayerCollapse introduces a theoretical regularizer and demonstrates significant efficiency gains across computer vision and natural language processing domains.
7 Conclusion and Limitations
LayerCollapse advances the development of efficient neural networks by optimizing performance alongside a reduction in model size and computational needs. However, the utility of LayerCollapse is somewhat restricted as it primarily compresses fully connected layers, which limits its effectiveness in CNN architectures. Moreover, this method necessitates either finetuning or pretraining to prepare the model for compression, and it also demands precise adjustment of the hyperparameter $\alpha$ to prevent a decline in performance and accuracy. These limitations highlight areas for further research to broaden the application of LayerCollapse to a more diverse array of network structures and model types.